40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 19 00:07:16 2013, Jenne, Update, LSC, Green status 
    Reply  Tue Nov 19 01:11:19 2013, Jenne, Update, LSC, Green status 
    Reply  Tue Nov 19 14:47:59 2013, manasa, Update, LSC, Green status 
       Reply  Tue Nov 19 17:47:17 2013, Jenne, Update, LSC, PRMI+2arms attempt 
Message ID: 9411     Entry time: Tue Nov 19 14:47:59 2013     In reply to: 9405     Reply to this: 9413
Author: manasa 
Type: Update 
Category: LSC 
Subject: Green status 

Quote:

After I aligned the IR interferometer (no ASS - we still need to figure out what's going on with that), I am trying to find the green beatnotes for each arm.

First, I locked the green lasers to each arm.

I then went out to the PSL table and aligned the Green Yarm path by overlapping the near-field and far-field of the yarm transmission and the PSL green pickoff.  I then turned on the power for the Beat PDs, since it was off (I confirmed that the outputs were plugged into the beatbox, so they are seeing 50 ohms).  I assume that the beat PDs were off since Manasa pulled the Beatbox last week, but there is no elog reference!!  Anyhow, after seeing a real signal, I maximized the DC power on the beat PD for the Yarm.  I then maximized the light on the DC transmission PD for the Yarm.

I looked at the Xarm, and the near-field alignment looks okay, but I haven't checked the far-field.

I started looking for the beatnotes from the control room:

I am changing the SLOW_SERVO2_OFFSETs by 30 counts, and then unlocking and relocking the arms, and checking to see if I see a peak on the RF spectrum analyser. 

The Y offset started at -10320, and I found a beatnote at -11230 (beatnote is about 26MHz).  The X offset started at 4500.  Going larger seemed to get me to a less bright TEM00 mode, so I switched and have been searching by going down in offset, but haven't yet found the beatnote.  I suspect that I actually need to align the X path on the PSL table.  The Y beatnote is very small, about -30dBm, so I also need to tweak the alignment by maximizing the peak value.

 I found the beatnotes for both the X and Y arm ALS this morning. The beat amplitudes measured -5dBm and -18dBm respectively and occurred at SLOW SERVO2 OFFSET 4550 and -10340. I had to only tweak the Y green PSL alignment to increase the beat amplitude.

I locked both the arms using ALS and they were stably locked until MC unlocked for a moment (nearly 16 minutes).

The only thing missing in the list of things you looked into is the status of the PSL slow actuator adjust. Check if this is near zero.

ELOG V3.1.3-